# Handling Missingness, Failures, and Non-Convergence in Simulation Studies: A Review of Current Practices and Recommendations

### v2

@@ 1 Introduction
This paper has two primary goals: First, we aim to systematically assess how the issue of missingness in simulation studies is currently reported and handled in the methodological research literature. Second, we aim to provide practical recommendations for methodological researchers on approaching missingness. To this end, we systematically reviewed research articles published in various methodological journals, summarized in Section 2. We then propose a classification of different types of missingness (Section 3) and outline approaches to dealing with them, including their strengths and limitations (Section 4). In Section 5, we use a case study on methods for publication bias adjustment in meta-analysis to illustrate how missingness can affect results and how to respond to it. Finally, we provide practical recommendations (Section 6) and conclude with a discussion of our findings, limitations, and implications (Section 7).
@@ 2.1 Sampling and coding procedure
Figure 1: Main results of the literature review of the Journal of the American Statistical Association (JASA), Psychological Methods (PM), Research Synthesis Methods (RSM), Statistics in Medicine (SiM). In panels A and C, we show per-journal percentages on the y-axis and proportions of articles with simulation studies and simulation studies with code, respectively, as fractions inside the bars. In panel B, “yes (zero)” refers to explicitly reporting that no missingness occurred. In panel C, “case-wise” and “list-wise” refer to the deletion of missing values, “add sims” refers to conducting additional simulation repetitions until a desired number is achieved, “replace” refers to the replacement of a method with another one. For panel D, we checked whether code for the simulation study and not just the implementation of the method was available. We did not check whether we could run the code.
@@ 2.4 Handling missingness
For the studies that elaborated on how missingness was handled when it occurred, we recorded the typical strategies of deletion (“case-wise” or “list-wise”), “additional simulation repetitions”, and “method replacement” (see panel C of Figure 1), each will be discussed in more detail in Section 4. This includes some studies in which the wording was not fully clear, but some handling was reported clearly (overall $n$ = 81). Hence, these studies include studies coded as “unclear” in panel B. Case-wise deletion denotes that only those per-method cases containing missing values were omitted in the performance evaluation. This was the most commonly applied strategy among studies that provided any details on missingness handling (used in 35/81 = 43% of the studies). Case-wise deletion was furthermore the most common strategy that was not explicitly reported but implicitly assumed by the coders (unquantified). List-wise deletion (9/81 = 11%) involves not only deleting the missing value itself, but also values from the other methods in the same simulation repetition. Only a small number of studies chose to perform additional simulation repetitions to compensate for the loss of deleted repetitions (4/81 = 5%) or replaced the missing values (4/81 = 5%), for example with the output of a different method. Other strategies (used in 29/81 = 36% of the studies) often combined two handling approaches. For example, Liu & Perera (2022) either performed additional simulation repetitions or did not interpret the performance of a method in case of a very high non-convergence rate.
@@ 4 Approaches for assessing and handling missingness
Case-wise deletion
@@
List-wise deletion
@@
Perform additional simulation repetitions to compensate for omitted repetitions, can be combined with case-wise or list-wise deletion or used to address DGM missingness
+ Different methods have the same desired number of repetitions, comparison based on the same data (when combined with list-wise deletion) Performance estimates might be non-representative of the intended DGM, computationally more intensive, hard to achieve with high missingness rates
@@ 4.1.1 Quantifying and reporting missingness
A good example is Castro-Alvarez et al. (2022), who differentiated between (i) errors and warnings, (ii) non-convergence, and (iii) resource limitations. They quantified the frequency of each per method and estimation approach in a table. Additionally, they visualized the number of successful analyses per condition. Huang et al. (2021) used a more minimal approach by adding a column to results tables showing converged cases – a simple practice that could become standard. Another example highlighting the importance of missingness quantification per method and condition is Hoyer & Kuss (2020). Individual conditions for one method had as low as 0.3% converged repetitions, casting doubt on its practical utility.
@@ 4.2.1 Non-analysis of conditions or methods with missingness
4.2.2 Case-wise and list-wise deletion of missing values
A simple strategy to estimate method performance in the presence of missingness is omitting the missing values via case-wise or list-wise deletion. However, both may implicitly lead to a different DGM than originally specified, since there is conditioning on non-missingness. This implicit DGM may not be representative of the specified DGM, especially if missingness is dependent on the characteristics of the simulated data sets (which it often is). To investigate this, it may be useful to look at descriptive statistics of the simulated data sets conditional on deletion. For example, if in a DGM a sample size is first simulated based on which a data set is then simulated, it may be useful to assess whether, within a condition, the average simulated sample size after deletion matches the expected sample size of the specified sample size distribution.
An additional problem with case-wise deletion is that if missingness varies between methods, their estimated performance is based on different underlying data, which can further distort the results. For example, if a particular method fails to converge on the most challenging data sets (e.g., data sets with small sample sizes), case-wise deletion may result in an over-optimistic assessment of its performance (Siepe et al., 2024; Wünsch et al., 2024). This is usually less of an issue with list-wise deletion since performance estimates are based on the same data for all methods, yet the implicit DGM underlying these performance estimates may be quite different from the specified DGM as there is a conditioning on simultaneous non-missingness of all methods. It is also essential to recognize that different missingness rates across methods are relevant for their practical utility. Finally, list-wise deletion also discards actual non-missing observations, thereby artificially increasing Monte Carlo uncertainty.
An interesting example from our literature review is the study by Seo et al. (2023) on multiple imputation in prediction modeling. These authors discuss the trade-off between case-wise and list-wise deletion, noting that list-wise deletion enables a fairer method comparison while case-wise deletion leads to much fewer omitted repetitions in conditions where multiple imputation does not converge. In the end, they decided on a combination of the two approaches, using list-wise deletion when the convergence rate of multiple imputation was below 20% in a condition, and case-wise deletion if it was above 20%.
@@ 4.2.3 Additional simulation repetitions
To avoid an increase in Monte Carlo uncertainty from omitting simulation repetitions with missing values, additional simulations can be run until the desired number of complete repetitions is reached (Paxton et al., 2001; Burton et al., 2006; Chalmers & Adkins, 2020). This can be computationally intensive and difficult to implement when missingness rates are high. In case of method missingness, the approach is typically combined with list-wise deletion. In principle, however, it could also be combined with case-wise deletion to avoid omitting non-missing repetitions, although this may lead to differing numbers of non-missing values among methods. When used to address DGM missingness, this approach resembles list-wise deletion as no performance is estimated for all methods for invalid data sets.
To give an example from our literature review, Lai & Hsiao (2022) used a list-wise deletion approach and performed additional simulation repetitions until they reached 5’000 repetitions where all compared methods converged to compensate for the omitted repetitions. However, this approach does not address the change in the implicit DGM, and could therefore still distort the results.
@@ 4.2.7 Method replacement
Several authors have suggested dealing with method missingness by replacing a missing value with a baseline or “gold standard” method (Crowther et al., 2014; Morris et al., 2019; White et al., 2024; Wünsch et al., 2024). This approach resembles how, on average, a method would be used in practice – if a method fails, a data analyst would not give up, but consider an alternative method or a method variant. Method replacement is therefore particularly useful for late-stage simulation studies evaluating the practical utility of specific method usage. However, choosing a replacement is not always straightforward. It also makes interpretation more difficult, since the performance estimates then refer to a mixture of two methods. For example, the empirical standard error of a method then relates to the combination of two methods, which may be quite different from the standard error estimate produced by the method in practice. Finally, a risk with this approach is that a method A may appear worse overall than a method B simply because it tends to be missing more often and its performance is worsened by the replacement method in the missing repetitions. Yet A might outperform B in conditions where both converge. To identify this, sensitivity analyses comparing case/list-wise deletion and method replacement approaches are necessary (see Section 6).
@@ 5 Case study
In our literature review, the simulation study would rank among the best-reported studies; Carter et al. (2019) provided per method-condition non-convergence rates, clearly stated that the performance measures were based on case-wise deletion, acknowledged method limitations due to non-convergence, and openly shared code and data for reproducing their results (https://osf.io/rf3ys/). In the following, we will use these data to reanalyze their simulation study.
@@ 5.2 Comparing missingness handling approaches
First, we consider the results under case-wise deletion (green) as reported by Carter et al. (2019). We see that all but the RE and p-curve methods show type I error rates below the nominal 5% across the three QRPs conditions. RE shows a slightly too high type I error rate in medium and high QRPs conditions, while the type I error rate of p-curve is slightly above 5% in the no QRPs condition (although 5% is included within the MCSE error bar). However, the interpretation of all these numbers is conditional on convergence, which drastically differs among the methods. As such, the comparison is not based on the same underlying data and might be unfairly biased because one method has more “difficult” or “easy” data sets to deal with than another.
@@
Second, in an effort to base method comparison on the same underlying data, one may omit all repetitions with at least one method showing non-convergence (list-wise deletion). From the blue statistics in Figure 5, we see that such an approach alters the results to some extent: First, it leads to larger MCSEs because more data are discarded than under the case-wise deletion approach. For instance, in the no QRPs condition (left panel), 79% of the repetitions are discarded, as only in 21% of all repetitions all methods converged. Furthermore, the estimated type I error rates of the methods can change drastically compared to case-wise deletion. For instance, in the no QRPs condition (left panel), the type I error rate of the RE method changes from about 2% (well below the nominal value of 5%) under case-wise deletion to about 7% under list-wise deletion. The choice between case-wise and list-wise deletion can thus determine whether a method is judged too conservative (below 5%) or too liberal (above 5%). While this comparison ensures that the comparison of method performance is based on the same underlying data, the interpretation is delicate (perhaps even more so than with case-wise deletion) because the estimated type I error rates are conditional on the simultaneous convergence of all compared methods and as such may not be particularly relevant in practice.
@@
Third, to mimic how a non-convergent method would actually be used in practice, we may also replace missing outputs from non-convergent methods with those from a “baseline” method. For example, we may use RE meta-analysis as a baseline method since it does not provide any adjustment for publication bias and always converges. Consequently, the approach leads to all methods having no missingness and thus the smallest MCSEs (see the orange statistics in Figure 5). At the same time, the interpretation becomes more difficult as the methods are no longer “pure”, e.g., the p-curve/uniform methods in the no QRPs condition (left panel) consist to 77% of the RE meta-analysis and only 23% of p-curve/uniform. As a result, the estimated type I error rate becomes closer to the type I error rate of RE in conditions with high missingness. For example, the p-curve method’s estimated type I error rate changes from around 6% under case-wise deletion to 2% under RE method replacement. While this approach keeps all repetitions and may be the most relevant for using a method in practice, it obscures the “true” method performance by replacing the method when non-convergence occurs.
@@ 5.3 Conclusions
Third, the comparison of the different handling approaches showed that for the investigated conditions, conclusions and interpretations were mostly consistent between case-/list-wise deletion and method replacement, but could differ drastically for certain methods and conditions based on the chosen handling approach. For example, the type I error rate of random effects meta-analysis changed from too conservative to too liberal in some conditions when switching from case-wise deletion or method replacement to list-wise deletion. Again, domain knowledge can be used to put these results into context. The type I error rate of random effects meta-analysis changed so radically because list-wise deletion removes all repetitions with only non-significant p-values, thus making the implicit DGM more difficult in the sense that more significant studies are observed even though the true effect is actually zero, and consequently worsening the performance of the method. With this in mind and given the goal of the study (to compare the performance under realistic conditions to give recommendations for application), list-wise deletion does not seem to be an appropriate handling approach. Baseline method replacement, on the other hand, seems to be a viable alternative to case-wise deletion chosen by the original authors, as it mimics how an applied researcher might use a publication bias adjustment method – if the publication bias method does not converge, revert to the unadjusted random effects meta-analysis baseline.
@@ 0. When observing missingness, reconsider methodology and implementation of method and simulation study
5. Perform sensitivity analyses to see how the results change when alternative missingness handling approaches are used (e.g., compare results of case-wise and list-wise deletion if both are justifiable)
@@ 6. Quantify and report missingness (ideally by method/condition, use supplement if space is limited)
When implementing a simulation study, it is common for initial attempts to produce issues that result in some form of missingness. In this case, the first step should be to rethink the implementation of the DGM, methods, and performance measures. In particular, when a new method is developed and missingness is encountered, this is often indicative of its limitations. Researchers should then reconsider the methodology and implementation of the method to avoid missingness. However, while iteratively adjusting a method or simulation study to understand and avoid missingness is a normal part of early method development, researchers should also be careful not to selectively choose favorable settings to present an unreliable method as better than it actually is (Pawel et al., 2024; Nießl et al., 2022).
@@
To explore the potential impact of the choice of the missingness handling approach, one may also perform sensitivity analyses to see how results change when alternative approaches are used (e.g., list-wise instead of case-wise deletion if both are justifiable). Such analyses become more important with increasing missingness. However, it is difficult to judge “how much missingness” is required until careful handling is necessary.

### v3

@@ 1 Introduction
This paper has two primary goals: First, we aim to systematically assess how the issue of missingness in simulation studies is currently reported and handled in the methodological research literature. Second, we aim to provide practical recommendations for methodological researchers on approaching missingness. To this end, we systematically reviewed research articles published in various methodological journals, summarized in Section 2. We then propose a classification of different types of missingness (Section 3) and outline approaches to dealing with them, including their strengths and limitations (Section 4). In Section 5, we use a case study on methods for publication bias adjustment in meta-analysis to illustrate how missingness can affect results and suggest ways of handling it. Finally, we provide practical recommendations (Section 6) and conclude with a discussion of our findings, limitations, and implications (Section 7).
@@ 2.1 Sampling and coding procedure
Figure 1: Main results of the literature review of the Journal of the American Statistical Association (JASA), Psychological Methods (PM), Research Synthesis Methods (RSM), Statistics in Medicine (SiM). In panels A and D, we show per-journal percentages on the y-axis and proportions of articles with simulation studies and simulation studies with code, respectively, as fractions inside the bars. Panel B shows how many articles mentioned missingness and, if so, whether they described how it was handled. “Yes (zero)” refers to articles that explicitly reported that no missingness occurred. In panel C, “method-wise” and “repetition-wise” refer to the deletion of missing values, “add sims” refers to conducting additional simulation repetitions until a desired number is achieved, “replace” refers to the replacement of a method with another one. For panel D, we checked whether code for the simulation study and not just the implementation of the method was available. We did not check whether we could run the code.
@@ 2.4 Handling missingness
For the studies that elaborated on how missingness was handled when it occurred, we recorded the typical strategies of deletion (“method-wise” or “repetition-wise”), “additional simulation repetitions”, and “method replacement” (see panel C of Figure 1), each will be discussed in more detail in Section 4. This includes some studies in which the wording was not fully clear, but some handling was reported clearly (overall $n$ = 81). Hence, these studies include studies coded as “unclear” in panel B. Method-wise deletion denotes that only those per-method cases containing missing values were omitted in the performance evaluation. This was the most commonly applied strategy among studies that provided any details on missingness handling (used in 35/81 = 43% of the studies). Method-wise deletion was furthermore the most common strategy that was not explicitly reported but implicitly assumed by the coders (unquantified). Repetition-wise deletion (9/81 = 11%) involves not only deleting the missing value itself, but also values from the other methods in the same simulation repetition. Only a small number of studies chose to perform additional simulation repetitions to compensate for the loss of deleted repetitions (4/81 = 5%) or replaced the missing values (4/81 = 5%), for example with the output of a different method. Other strategies (used in 29/81 = 36% of the studies) often combined two handling approaches. For example, Liu & Perera (2022) either performed additional simulation repetitions or did not interpret the performance of a method in case of a very high non-convergence rate.
@@ 4 Approaches for assessing and handling missingness
Method-wise deletion
@@
Repetition-wise deletion
@@
Perform additional simulation repetitions to compensate for omitted repetitions, can be combined with method-wise or repetition-wise deletion or used to address DGM missingness
+ Different methods have the same desired number of repetitions, comparison based on the same data (when combined with repetition-wise deletion) Performance estimates might be non-representative of the intended DGM, computationally more intensive, hard to achieve with high missingness rates
@@ 4.1.1 Quantifying and reporting missingness
A good example is Castro-Alvarez et al. (2022), who differentiated between (i) errors and warnings, (ii) non-convergence, and (iii) resource limitations. They quantified the frequency of each per method and estimation approach in a table. Additionally, they visualized the number of successful analyses per condition. Huang et al. (2021) used a more minimal approach by adding a column to results tables showing the number of converged cases – a simple practice that could become standard. Another example highlighting the importance of missingness quantification per method and condition is Hoyer & Kuss (2020). Individual conditions for one method had as low as 0.3% converged repetitions, casting doubt on its practical utility.
@@ 4.2.1 Non-analysis of conditions or methods with missingness
4.2.2 Method-wise and repetition-wise deletion of missing values
A simple strategy to estimate method performance in the presence of missingness is omitting the missing values via method-wise or repetition-wise deletion. However, both may implicitly lead to a different DGM than originally specified, since there is conditioning on non-missingness. This implicit DGM may not be representative of the specified DGM, especially if missingness is dependent on the characteristics of the simulated data sets (which it often is). To investigate this, it may be useful to look at descriptive statistics of the simulated data sets conditional on deletion. For example, if in a DGM a sample size is first simulated based on which a data set is then simulated, it may be useful to assess whether, within a condition, the average simulated sample size after deletion matches the expected sample size of the specified sample size distribution.
An additional problem with method-wise deletion is that if missingness varies between methods, their estimated performance is based on different underlying data. This can make it difficult to compare the precision of performance estimates (e.g. with Monte Carlo standard errors) as a different number of repetitions may be available for each method. It can further distort the results. For example, if a particular method fails to converge on the most challenging data sets (e.g., data sets with small sample sizes), method-wise deletion may result in an over-optimistic assessment of its performance (Siepe et al., 2024; Wünsch et al., 2024). This is usually less of an issue with repetition-wise deletion since performance estimates are based on the same data for all methods, yet the implicit DGM underlying these performance estimates may be quite different from the specified DGM as there is a conditioning on simultaneous non-missingness of all methods. It is also essential to recognize that different missingness rates across methods are relevant for their practical utility. Finally, repetition-wise deletion also discards actual non-missing observations, thereby artificially increasing Monte Carlo uncertainty.
An interesting example from our literature review is the study by Seo et al. (2023) on multiple imputation in prediction modeling. These authors discuss the trade-off between method-wise and repetition-wise deletion, noting that repetition-wise deletion enables a fairer method comparison while method-wise deletion leads to much fewer omitted repetitions in conditions where multiple imputation does not converge. In the end, they decided on a combination of the two approaches, using repetition-wise deletion when the convergence rate of multiple imputation was below 20% in a condition, and method-wise deletion if it was above 20%.
@@ 4.2.3 Additional simulation repetitions
To avoid an increase in Monte Carlo uncertainty from omitting simulation repetitions with missing values, additional simulations can be run until the desired number of complete repetitions is reached (Paxton et al., 2001; Burton et al., 2006; Chalmers & Adkins, 2020). This can be computationally intensive and difficult to implement when missingness rates are high. In case of method missingness, the approach is typically combined with repetition-wise deletion. However, in principle, it could also be combined with method-wise deletion to avoid omitting non-missing values. That is, one could still retain repetitions where some methods are missing and perform additional simulations until all methods have at least the desired number of non-missing values. This may result in different numbers of non-missing values among methods, as some methods may have more non-missing values than the desired number of repetitions. Finally, when used to address DGM missingness, the additional simulation approach resembles repetition-wise deletion as no performance is estimated for all methods for invalid data sets.
To give an example from our literature review, Lai & Hsiao (2022) used a repetition-wise deletion approach and performed additional simulation repetitions until they reached 5’000 repetitions where all compared methods converged to compensate for the omitted repetitions. However, this approach does not address the change in the implicit DGM, and could therefore still distort the results.
@@ 4.2.7 Method replacement
Several authors have suggested dealing with method missingness by replacing a missing value with a baseline or “gold standard” method (Crowther et al., 2014; Morris et al., 2019; White et al., 2024; Wünsch et al., 2024). This approach resembles how, on average, a method would be used in practice – if a method fails, a data analyst would not give up, but consider an alternative method or a method variant. Method replacement is therefore particularly useful for late-stage simulation studies evaluating the practical utility of specific method usage. However, choosing a replacement is not always straightforward. It also makes interpretation more difficult, since the performance estimates then refer to a mixture of two methods. For example, the empirical standard error of a method then relates to the combination of two methods, which may be quite different from the standard error estimate produced by the method in practice. Finally, a risk with this approach is that a method A may appear worse overall than a method B simply because it tends to be missing more often and its performance is worsened by the replacement method in the missing repetitions. Yet A might outperform B in conditions where both converge. To identify this, sensitivity analyses comparing method-/repetition-wise deletion and method replacement approaches are necessary (see Section 6).
@@ 5 Case study
In our literature review, the simulation study would rank among the best-reported studies; Carter et al. (2019) provided per method-condition non-convergence rates, clearly stated that the performance measures were based on method-wise deletion, acknowledged method limitations due to non-convergence, and openly shared code and data for reproducing their results (https://osf.io/rf3ys/). In the following, we will use these data to reanalyze their simulation study.
@@ 5.2 Comparing missingness handling approaches
First, we consider the results under method-wise deletion (green) as reported by Carter et al. (2019). We see that all but the RE and p-curve methods show type I error rates below the nominal 5% across the three QRPs conditions. RE shows a slightly too high type I error rate in medium and high QRPs conditions, while the type I error rate of p-curve is slightly above 5% in the no QRPs condition (although 5% is included within the MCSE error bar). However, the interpretation of all these numbers is conditional on convergence, which drastically differs among the methods. As such, the comparison is not based on the same underlying data and might be unfairly biased because one method has more “difficult” or “easy” data sets to deal with than another.
@@
Second, in an effort to base method comparison on the same underlying data, one may omit all repetitions with at least one method showing non-convergence (repetition-wise deletion). From the blue statistics in Figure 5, we see that such an approach alters the results to some extent: First, it leads to larger MCSEs because more data are discarded than under the method-wise deletion approach. For instance, in the no QRPs condition (left panel), 79% of the repetitions are discarded, as only in 21% of all repetitions all methods converged. Furthermore, the estimated type I error rates of the methods can change drastically compared to method-wise deletion. For instance, in the no QRPs condition (left panel), the type I error rate of the RE method changes from about 2% (well below the nominal value of 5%) under method-wise deletion to about 7% under repetition-wise deletion. The choice between method-wise and repetition-wise deletion can thus determine whether a method is judged too conservative (below 5%) or too liberal (above 5%). While this comparison ensures that the comparison of method performance is based on the same underlying data, the interpretation is delicate (perhaps even more so than with method-wise deletion) because the estimated type I error rates are conditional on the simultaneous convergence of all compared methods and as such may not be particularly relevant in practice.
@@
Third, to mimic how a non-convergent method would actually be used in practice, we may also replace missing outputs from non-convergent methods with those from a “baseline” method. For example, we may use RE meta-analysis as a baseline method since it does not provide any adjustment for publication bias and always converges. Consequently, the approach leads to all methods having no missingness and thus the smallest MCSEs (see the orange statistics in Figure 5). At the same time, the interpretation becomes more difficult as the methods are no longer “pure”, e.g., the p-curve/uniform methods in the no QRPs condition (left panel) consist to 77% of the RE meta-analysis and only 23% of p-curve/uniform. As a result, the estimated type I error rate becomes closer to the type I error rate of RE in conditions with high missingness. For example, the p-curve method’s estimated type I error rate changes from around 6% under method-wise deletion to 2% under RE method replacement. While this approach keeps all repetitions and may be the most relevant for using a method in practice, it obscures the “true” method performance by replacing the method when non-convergence occurs.
@@ 5.3 Conclusions
Third, the comparison of the different handling approaches showed that for the investigated conditions, conclusions and interpretations were mostly consistent between method-/repetition-wise deletion and method replacement, but could differ drastically for certain methods and conditions based on the chosen handling approach. For example, the type I error rate of random effects meta-analysis changed from too conservative to too liberal in some conditions when switching from method-wise deletion or method replacement to repetition-wise deletion. Again, domain knowledge can be used to put these results into context. The type I error rate of random effects meta-analysis changed so radically because repetition-wise deletion removes all repetitions with only non-significant p-values, thus making the implicit DGM more difficult in the sense that more significant studies are observed even though the true effect is actually zero, and consequently worsening the performance of the method. With this in mind and given the goal of the study (to compare the performance under realistic conditions to give recommendations for application), repetition-wise deletion does not seem to be an appropriate handling approach. Baseline method replacement, on the other hand, seems to be a viable alternative to method-wise deletion chosen by the original authors, as it mimics how an applied researcher might use a publication bias adjustment method – if the publication bias method does not converge, revert to the unadjusted random effects meta-analysis baseline.
@@ 0. When observing missingness, reconsider methodology and implementation of method and simulation study
5. Perform sensitivity analyses to see how the results change when alternative missingness handling approaches are used (e.g., compare results of method-wise and repetition-wise deletion if both are justifiable)
@@ 6. Quantify and report missingness (ideally by method/condition, use supplement if space is limited)
When implementing a simulation study, it is common for initial attempts to produce issues that result in some form of missingness. In this case, the first step should be to rethink the implementation of the DGM, methods, and performance measures. In particular, when a new method is developed and missingness is encountered, this is often indicative of its limitations. Researchers should then reconsider the methodology and implementation of the method to avoid missingness before even attempting to address it in any of the ways discussed in this paper (therefore indicated as “recommendation 0” in Table 3). However, while iteratively adjusting a method or simulation study to understand and avoid missingness is a normal part of early method development, researchers should also be careful not to selectively choose favorable settings to present an unreliable method as better than it actually is (Pawel et al., 2024; Nießl et al., 2022).
@@
To explore the potential impact of the choice of the missingness handling approach, one may also perform sensitivity analyses to see how results change when alternative approaches are used (e.g., repetition-wise instead of method-wise deletion if both are justifiable). Such analyses become more important with increasing missingness. However, it is difficult to judge “how much missingness” is required until careful handling is necessary. Sensitivity analyses can also be useful for providing more differentiated method recommendations. For example, two result tables could be presented: one showing performance estimates based on method-wise deletion and another showing estimates based on gold standard replacement. This could then inform two-part recommendations, such as using a certain method when convergence occurs (when it performs best) but using another method when convergence does not occur.